A randomly selected sample of college baseball players has the following heights in inches. 68, 63, 66, 63, 68, 63, 65, 66, 65, 67, 65, 65, 69, 71, 65, 70, 61, 66, 69, 62, 65, 64, 70, 63, 71, 63, 68, 68, 62, 71, 62, 65 Compute a 95% confidence interval for the population mean height of college baseball players based on this sample and fill in the blanks appropriately. < μ < (Keep 3 decimal places)
Answer #1

solution Here from the given data n=32 ă=65.906, s=2.955 d-f=n-1=32-1=31 At 95% confidence taj - 2.04 The 95% c-I for I is to